# LiNeS: Post-training Layer Scaling Prevents Forgetting and Enhances Model Merging

### v1

@@ 5.2 Improving Multi-task Model Merging
In this section, we extend LiNeS to improve multi-task merging algorithms, aiming to combine multiple models fine-tuned independently on different tasks into a single model (Matena & Raffel, 2022; Ilharco et al., 2023; Ortiz-Jimenez et al., 2023; Yadav et al., 2023). Task arithmetic (Ilharco et al., 2023) proposed to decouple the contributions of the pre-trained model and individual task vectors, first generating a multi-task vector $\bm{\tau}_{\textrm{MTL}}=g(\bm{\tau}_{1},\dots,\bm{\tau}_{T})$ with a merging function $g:\mathbb{R}^{N}\times\cdots\times\mathbb{R}^{N}\mapsto\mathbb{R}^{N}$, and then adding back to the pre-trained checkpoint with a scaling factor to construct a multi-task model $\bm{\theta}=\bm{\theta}_{0}+\lambda\cdot\bm{\tau}_{\textrm{MTL}}$. The scalar coefficient $\lambda$ is tuned using a held-out validation set. Recent works (Yadav et al., 2023; Wang et al., 2024) follow the same protocol while improving the merging function $g$ for retaining more task information. We refer to Appendix B.1 for a more detailed explanation of these methods.
$\alpha=\frac{1}{N_{\textrm{models}}}\frac{\left\|\bm{\tau}_{\text{sum}}\right\|}{\|\bm{\tau}_{\textrm{MTL}}\|},\textrm{ where }\tau_{\text{sum}}=\sum_{i=1}^{N_{\textrm{models}}}{\bm{\tau}_{i}}$ (2)
Therefore, we only tune $\beta$ and search over the same range as the constant scaling $\lambda$ used by the aforementioned merging techniques. As a result, LiNeS shares the same computational requirements as the baseline merging methods; we provide more details for the hyper-parameters in Appendix B.2.1. Specifically, we consider various multi-task model merging baselines. namely Task Arithmetic (Ilharco et al., 2023), Ties-merging (Yadav et al., 2023), Consensus Merging (Wang et al., 2024), enhancing them with LiNeS and evaluate on both computer vision and NLP benchmarks.
@@ 5.3 Improving Model Soups for Merging single-task models
Averaging in weight space multiple models fine-tuned on the same task derived from the same pre-trained model has been shown to increase target performance (Wortsman et al., 2022a; Rame et al., 2022). In this section, we investigate whether LiNeS can enhance the test performance when merging single-task models.
@@ 5.4 Improving Rewarded Soups
Starting with an LLM parameterized by weights $\bm{\theta}_{0}$, we first fine-tune it using supervised fine-tuning (SFT) on labeled demonstrations. From the resulting weights $\bm{\theta}_{\rm SFT}$, we then apply Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022), training two independent policies via Proximal Policy Optimization (PPO) (Schulman et al., 2017) to maximize the rewards $R_{1}$ and $R_{2}$ respectively. To merge these policies, we linearly interpolate the residuals $\bm{\tau}_{1}=\bm{\theta}_{1}-\bm{\theta}_{\rm SFT}$ and $\bm{\tau}_{2}=\bm{\theta}_{2}-\bm{\theta}_{\rm SFT}$. This interpolation defines a continuous set of rewarded policies:
@@
In our experiment, we use LLaMA-2 7B (Touvron et al., 2023) and the Reddit Summary task (Stiennon et al., 2020), which consists of 14.9k post-summary pairs. We employ two reward models: GPT2-reward-summarization – which scores summaries based on human preferences – and BART-faithful-summary-detector (Chen et al., 2021) – which evaluates the faithfulness of the generated summary to the source post. To evaluate the models, we use a subset of 1k samples from the test set, generate the responses, and compute the average score for each reward dimension.
@@ 7 Conclusion
In this work, we presented LiNeS, a novel method designed to mitigate catastrophic forgetting after fine-tuning process. By reducing the magnitude for parameter updates in the shallower layers, LiNeS improves the generalization performance of the edited model on control tasks while almost fully preserving performance on the fine-tuned tasks. Furthermore, we demonstrated the versatility of LiNeS in addressing task interference in multi-task model merging, where it consistently improves the baseline model merging methods across vision and NLP benchmarks. Our experiments confirmed the broad applicability of LiNeS across various scenarios, from improving OOD generalization to enhancing multi-task and single-task model merging strategies, as well as improving merging LLM policies aligned with different rewards. Given its simplicity and ease of integration with existing methods, LiNeS offers a practical and inexpensive solution for boosting the generalization and robustness of fine-tuned models in diverse application domains.
@@ Acknowledgments
The authors thank Adam Hazimeh for constructive discussions and comments.
@@ Appendix A LiNeS Pseudocode
Parameters: ———– task_vector : dict A dictionary representing the residual between the fine-tuned checkpoint and the pre-trained checkpoint. alpha : float The minimum scaling factor for the blocks. beta : float The maximum scaling coefficient difference between the last and first block. num_blocks : int The total number of layer blocks in the model. Returns: ——– scaled_task_vector : dict A copy of ‘task_vector‘ where each key is scaled based on the layer depth. ”””
@@
# Create a scaling dictionary to store the scaling factor for each key scaling_dic = for k in task_vector.keys(): # Find the layer block in the key and assign scaling factor based on layer depth for layer, block in enumerate(key_blocks): if block in k: scaling_dic[k] = alpha + beta * (layer / (num_blocks - 1)) break
@@ B.2.2 Benchmarks
  • The 14-task benchmark includes the original eight tasks plus additional ones: CIFAR100 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), Flowers102 (Nilsback & Zisserman, 2008), OxfordIIITPet (Parkhi et al., 2012), PCAM (Veeling et al., 2018), and FER2013 (Goodfellow et al., 2013).
  • The 20-task benchmark builds on the 14-task benchmark with the addition of: EMNIST (Cohen et al., 2017), CIFAR10 (Krizhevsky et al., 2009), Food101 (Bossard et al., 2014), FashionMNIST (Xiao et al., 2017), RenderedSST2 (Socher et al., 2013; Radford et al., 2019), and KMNIST (Clanuwat et al., 2018).
@@ C.5 Results for editing T5 with LiNeS
C.6 Additional results for improving WiSE-FT with LiNeS
C.6.1 Results for using ViT-B/16 as visual encoder
@@
C.6.2 Individual results for 70 checkpoints
@@
C.7 Detailed performance on individual tasks for multi-task model merging

### v2

@@ 5.2 Improving Multi-task Model Merging
In this section, we extend LiNeS to improve multi-task merging algorithms, aiming to combine multiple models fine-tuned independently on different tasks into a single model (Matena & Raffel, 2022; Ilharco et al., 2023; Ortiz-Jimenez et al., 2023; Yadav et al., 2023; Hazimeh et al., 2024). Task arithmetic (Ilharco et al., 2023) proposed to decouple the contributions of the pre-trained model and individual task vectors, first generating a multi-task vector $\bm{\tau}_{\textrm{MTL}}=g(\bm{\tau}_{1},\dots,\bm{\tau}_{T})$ with a merging function $g:\mathbb{R}^{N}\times\cdots\times\mathbb{R}^{N}\mapsto\mathbb{R}^{N}$, and then adding back to the pre-trained checkpoint with a scaling factor to construct a multi-task model $\bm{\theta}=\bm{\theta}_{0}+\lambda\cdot\bm{\tau}_{\textrm{MTL}}$. The scalar coefficient $\lambda$ is tuned using a held-out validation set. Recent works (Yadav et al., 2023; Wang et al., 2024) follow the same protocol while improving the merging function $g$ for retaining more task information. We refer to Appendix B.1 for a more detailed explanation of these methods.
$\alpha=\frac{1}{N_{\textrm{models}}}\frac{\left\|\bm{\tau}_{\text{sum}}\right\|}{\|\bm{\tau}_{\textrm{MTL}}\|},\textrm{ where }\bm{\tau}_{\text{sum}}=\sum_{i=1}^{N_{\textrm{models}}}{\bm{\tau}_{i}}$ (2)
Therefore, we only tune $\beta$ and search over the same range as the constant scaling $\lambda$ used by the aforementioned merging techniques. As a result, LiNeS shares the same computational requirements as the baseline merging methods; we provide more details for the hyper-parameters in Appendix B.2.1, as well as a sensitivity analysis to the hyper-parameters in Appendix C.6. Specifically, we consider various multi-task model merging baselines, namely Task Arithmetic (Ilharco et al., 2023), Ties-merging (Yadav et al., 2023), Consensus Merging (Wang et al., 2024), enhancing them with LiNeS and evaluate on both computer vision and NLP benchmarks.
@@ 5.3 Improving Model Soups for Merging single-task models
Averaging in weight space multiple models fine-tuned on the same task derived from the same pre-trained model has been shown to increase target performance (Wortsman et al., 2022a; Ramé et al., 2022). In this section, we investigate whether LiNeS can enhance the test performance when merging single-task models.
@@ 5.4 Improving Rewarded Soups
Starting with an LLM parameterized by weights $\bm{\theta}_{0}$, we first fine-tune it using supervised fine-tuning (SFT) on labeled demonstrations. From the resulting weights $\bm{\theta}_{\rm SFT}$, we then apply Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022), training two independent policies via Proximal Policy Optimization (PPO) (Schulman et al., 2017) to maximize the rewards $R_{1}$ and $R_{2}$ respectively. To merge these policies, we linearly interpolate the residuals $\bm{\tau}_{1}=\bm{\theta}_{1}-\bm{\theta}_{\rm SFT}$ and $\bm{\tau}_{2}=\bm{\theta}_{2}-\bm{\theta}_{\rm SFT}$, defining a continuous set of rewarded policies:
@@
In our experiment, we use LLaMA-2 7B (Touvron et al., 2023) and the Reddit Summary task (Stiennon et al., 2020), which consists of 14.9k post-summary pairs. We fine-tune the model using LoRA (Hu et al., 2022) with $r_{\rm\scriptscriptstyle LoRA}=64$, $\alpha_{\rm\scriptscriptstyle LoRA}=128$, and 0.05 dropout. We employ two reward models: GPT2-reward-summarization – which scores summaries based on human preferences – and BART-faithful-summary-detector (Chen et al., 2021) – which evaluates the faithfulness of the generated summary to the source post. To evaluate the models, we use a subset of 1k samples from the test set, generate the responses, and compute the average score for each reward dimension.
@@ 7 Conclusion
In this work, we presented LiNeS, a novel method designed to mitigate catastrophic forgetting after fine-tuning process. By reducing the magnitude for parameter updates in the shallower layers, LiNeS improves the generalization performance of the edited model on control tasks while almost fully preserving performance on the fine-tuned tasks. Furthermore, we demonstrated the versatility of LiNeS in addressing task interference in multi-task model merging, where it consistently improves the baseline model merging methods across vision and NLP benchmarks. Our experiments confirm the broad applicability of LiNeS across various scenarios, from improving OOD generalization to enhancing multi-task and single-task model merging strategies, as well as improving merging LLM policies aligned with different rewards. Given its simplicity and ease of integration with existing methods, LiNeS offers a practical and inexpensive solution for boosting the generalization and robustness of fine-tuned models in diverse application domains.
@@ Acknowledgments
The authors thank Adam Hazimeh and the anonymous reviewers for their constructive discussions and comments.
@@ Appendix A LiNeS Pseudocode
Parameters: ———– task_vector : dict A dictionaryves control performantween the fine-tuned checkpoint and the pre-trained checkpoint. alpha : float The minimum scaling factor for the blocks. beta : float The maximum scaling coefficient difference between the last and first block. num_blocks : int The total number of layer blocks in the model. Returns: ——– scaled_task_vector : dict A copy of ‘task_vector‘ where each key is scaled based on the layer depth. ”””
@@
# Create a scaling dictionary to store the scaling factor for each key scaling_dic = for k in task_vector.keys(): # Find the layer block in the key and assign scaling factor based # on layer depth for layer, block in enumerate(key_blocks): if block in k: scaling_dic[k] = alpha + beta * (layer / (num_blocks - 1)) break
@@ B.2.2 Benchmarks
  • The 14-task benchmark includes the original eight tasks plus additional ones: CIFAR100 (Krizhevsky & Hinton, 2009), STL10 (Coates et al., 2011), Flowers102 (Nilsback & Zisserman, 2008), OxfordIIITPet (Parkhi et al., 2012), PCAM (Veeling et al., 2018), and FER2013 (Goodfellow et al., 2013).
  • The 20-task benchmark builds on the 14-task benchmark with the addition of: EMNIST (Cohen et al., 2017), CIFAR10 (Krizhevsky & Hinton, 2009), Food101 (Bossard et al., 2014), FashionMNIST (Xiao et al., 2017), RenderedSST2 (Socher et al., 2013; Radford et al., 2019), and KMNIST (Clanuwat et al., 2018).
@@ C.5 Results for editing T5 with LiNeS
C.6 Sensitivity analysis for hyper-parameters
We provide in this section the sensitivity analysis for the hyper-parameters of LiNeS. Specifically, we consider the setting in multi-task merging in the 8-task vision classification benchmark with ViT-B/32 CLIP model.
As explained in Section 5.2, LiNeS fixes $\alpha$ with a heuristic value by Equation 2 and only tunes $\beta$ for multi-task merging. The slope hyper-parameter $\beta$ is tuned within the same range as the uniform scaling coefficient $\lambda$ for the baseline merging methods. We compare in Figure 8 the sensitivity of averaged multi-task validation accuracy to the respective hyper-parameters, i.e., to $\lambda$ for baseline merging methods and $\beta$ for the LiNeS-enhanced merging methods. The results show that, for all three merging methods, including Task arithmetic, Ties-merging and Consensus, enhancing with LiNeS is less sensitive to hyper-parameter choices compared to the corresponding baseline method.
Figure 8: Sensitivity to hyper-parameters in multi-task merging for the 8-task benchmark with CLIP ViT-B/32 model. The y-axis represents the averaged multi-task validation accuracy and x-axis represents the hyper-parameter value, i.e., $\lambda$ for the baseline method and $\beta$ for the method enhanced with LiNeS.
Furthermore, we perform an ablation treating $\alpha$ as a hyper-parameter and analyze the sensitivity to both $\alpha$ and $\beta$ for LiNeS in a two-dimensional grid for the same benchmark. The results are presented in Figure 9. The results clearly demonstrate the necessity for applying layer-increasing scaling, as the optimal performance is obtained with both $\alpha>0$ and $\beta>0$ for all three merging method. Note that the optimal configurations found by the ablation study are very close to the configurations found in our method, as shown in Table 8, by setting $\alpha$ via the heuristic and searching only for $\beta$.
Figure 9: Sensitivity to both $\alpha$ and $\beta$ in multi-task merging for the 8-task benchmark with CLIP ViT-B/32 model. The heatmap represents the averaged multi-task validation accuracy, while x and y axis represent the $\beta$ and $\alpha$ respectively. The optimal configuration is annotated with a red box.
Table 8: $\alpha$ and $\beta$ values for $\alpha$ set by our proposed heuristic in Equation 2.
Method $\alpha$ $\beta$
Task Arithmetic 0.125 0.3
Ties-merging 0.21 1.4
Consensus merging 0.21 0.9
C.7 Experiments with CNN architectures
In this section, we apply LiNeS to CNN architectures. Specifically, we consider the ConvNeXt (Liu et al., 2022) architecture. First, we repeat the experiments presented in Section 3 regarding mitigating catastrophic forgetting. The final results are presented in Figure 10, where we observe similar findings with CLIP ViTs, i.e., LiNeS greatly improves the performance on control tasks when applied to the fine-tuned checkpoints while preserving most of the accuracy on target tasks. Furthermore, we present in Table 9 the results on multi-task model merging, following the experimental protocol established in Section 5.2. Again, we see that LiNeS improves the performance of baseline merging methods.
Figure 10: Our linear scaling (LiNeS) retains performance on both control and fine-tuned target tasks for ConvNext architecture.
Table 9: Multi-task model merging results using a ConvNeXt architecture. LiNeS improves the results compared to uniform scaling for both Task Arithmetic and Ties-merging.
Method LiNeS Acc (%) Norm. Acc (%)
Task Arithmetic 77.9 83.8
79.0 [+1.1] 84.8 [+1.0]
Ties-merging 79.7 85.8
80.3 [+0.6] 86.3 [+0.5]
C.8 Experiments with Regularized Fine-Tuning
In this section, we evaluate LiNeS against several regularized fine-tuning methods, focusing on their ability to preserve general features and mitigate catastrophic forgetting. The regularization strategies applied during fine-tuning are described below:
  1. 1.
    Fine-tuning with Linear Layer-Wise Learning Rate Decay (LinLR): Applies a linear learning rate schedule where the learning rate linearly increases from 0.0 to the maximum value for all the layers.
  2. 2.
    Fine-tuning with Exponential Layer-Wise Learning Rate Decay (ExpLR): Applies an exponential learning rate schedule where the learning rate is set to maximum for the deepest layers and decays by a factor of 0.5 by each layer for the shallower layers.
  3. 3.
    Fine-tuning with First Half of Blocks Frozen (HalfFT): Freezes the parameters of the first half of the model’s blocks during training.
  4. 4.
    Fine-tuning only the Final Block (LastFT): Freezes all blocks except the final block of the feature encoder.
Table 10: Performance of different methods on target and control tasks, averaged over all target and control task combinations in the 8-task vision benchmark (Ilharco et al., 2023).
pre-trained fine-tuned FT+LiNeS FT+LinLR FT+ExpLR FT+HalfFT FT+LastFT
Target (%) 48.3 90.5 90.3 90.7 89.6 90.4 85.6
Control (%) 48.3 38.0 48.0 46.9 46.0 46.8 46.6
We present the results in LABEL:table:regularized_ft. We observe that LiNeS, as a post-training editing method, outperforms the regularized fine-tuning methods in terms of restoring the zero-shot performance on the control tasks. We further emphasize that compared with the regularized fine-tuning methods, LiNeS benefits from many advantages such as efficiency, flexibility and computational cost.
C.9 Experiments with LoRA Fine-tuning
We explore the applicability of the method on models fine-tuned with LoRA (Hu et al., 2022). For a layer with pre-trained weights $\bm{W}_{0}\in\mathbb{R}^{m\times n}$, LoRA adds trainable matrices $\bm{A}\in\mathbb{R}^{m\times r}$ and $\bm{B}\in\mathbb{R}^{n}$, for rank $r\ll\min(n,m)$. The weights of the layer become:
$\bm{W}=\bm{W}_{0}+\frac{\alpha}{r}\bm{B}\bm{A}$
for $\alpha\in\mathbb{R}$. Following common practice, we set $\alpha=r$ and fine-tune with the same protocol used for full fine-tuning. We consider only the case of ViT-B/32 fine-tuned on 8 tasks and replicate the experiment presented in Table 1. Specifically, for each of the 8 LoRA-fine-tuned models, we compute the accuracy on the same (target) task as well as the average performance for each of the 7 remaining control tasks. Table 11 reports the average over the 8 cases for ranks $r\in\{16,32,64,128\}$. We observe that LoRA fine-tuning has lower target performance compared to full fine-tuning and that increasing target performance comes at the cost of more forgetting. In all the cases, LiNeS restores control performance while minimally affecting target performance.
Table 11: Similar to Table 1, LoRA Fine-tuning harms generalization on control tasks. Increased target performance rsults in higher levels of forgetting. Still, our proposed method LiNeS restores control performance for all ranks considered while minimally affecting target performance.
Target Control
Pre-trained 48.3 48.3
Fine-tuned 90.5 38.0
[…]
$r=16$ 84.4 44.2
$r=16+\texttt{LiNeS}$ 84.3 46.7
$r=32$ 85.8 42.8
$r=32+\texttt{LiNeS}$ 85.5 46.7
$r=64$ 86.6 41.6
$r=64+\texttt{LiNeS}$ 86.4 46.2
$r=128$ 87.5 41.6
$r=128+\texttt{LiNeS}$ 87.2 46.3
C.10 Additional results for improving WiSE-FT with LiNeS
C.10.1 Results for using ViT-B/16 as visual encoder
@@
C.10.2 Individual results for 70 checkpoints
@@
C.11 Detailed performance on individual tasks for multi-task model merging